OpenStudy (anonymous):

C3H2N, 208.17g/mol molecular formula?

OpenStudy (anonymous):

First you need to find the molecular mass of C3H2N. C*3 = 36.03g +H*2 = 2.02g +N = 14.01g = 52.05g/mol Now you just divide the molecular mass of the molecular formula, which is given, by the molar mass which we just found. \[\frac{208.17g/mol}{52.05g/mol} \approx 4\] Now you just multiply each of the subscripts of the empirical formula by the 4 we just got. C3H2N * 4 = C12H8N4 C12H8N4 would be the molecular formula.